Identify a pattern and find the next three numbers in the pattern.
1,3,9,27,81....

Found 2 solutions by jim_thompson5910, stanbon:
Answer by jim_thompson5910(35256)

Each previous term is being multiplied by 3 to get the next term

1*3 = 3
3*3 = 9
9*3 = 27
27*3 = 81
etc etc

The next three numbers would be

81*3 = 243
243*3 = 729
729*3 = 2187

Answer by stanbon(75887)

Identify a pattern and find the next three numbers in the pattern.
1,3,9,27,81, 3^5 = 243, 3^6 = 729, 3^7 = 2187....
-------
Cheers,
Stan H.
-------------